HOMOGENEOUS EQUATION

Solve the given equation below. Show complete solution.


xy' = x + y


1
Expert's answer
2021-06-07T18:43:13-0400

Ans:-

"xy'=x+y"


"\dfrac{dy}{dx}=\dfrac{x+y}{x}\ \ \ \ \ \ \ \ \ \ \ ...1"


Solving "\dfrac{dy}{dx}" by putting "\ \ y=vx"


differentiating w. r. t. x


"\dfrac{dy}{dx}=x\dfrac{dv}{dx}+v\dfrac{dx}{dx}"


"\dfrac{dy}{dx}=x\dfrac{dv}{dx}+v"


Putting value of "\dfrac{dy}{dx}" and y=vx in (1)


"\dfrac{dy}{dx}=\dfrac{x+y}{x}"


"x\dfrac{dv}{dx}+v=\dfrac{x+vx}{x}\\"


"x\dfrac{dv}{dx}+v=1+v\\"


"\dfrac{dv}{dx}=\dfrac{1}{x}"


Integrating both sides


"\int dv=\int \dfrac{dx}{x}"


"v=log|x| +c"


putting "v=\dfrac{y}{x}"


"\dfrac{y}{x}=log|x| +c"


"y=xlog|x|+cx"